A body of mass 20kg weights 34N on the moon. Find the mass of another body which weights 5100N under the same conditions.
1
Expert's answer
2020-06-15T10:22:47-0400

First, fond the acceleration due to gravity "g" on the moon:


"g=\frac{P_1}{m_1}."

The mass of another body is


"m_2=\frac{P_2}{g}=m_1 \frac{P_2}{P_1}=20 \cdot\frac{5100}{34}=3000\text{ kg}."